2Sin a=2-Cos A, FIND SinA
 

Grade:10

1 Answers

Sher Mohammad IIT Delhi
9 years ago
Cos A=2-2 Sina
SinA= sqrt9(1-Cos A 2)
=(1- 4(1-Sina)2)